OpenStudy (anonymous):

Consider the function f(x) = 6sin(x^2) on the interval 0 ≤ x ≤ 3. (a) Find the exact value in the given interval where an antiderivative, F, reaches its maximum. x = (b) If F(1) = 5, estimate the maximum value attained by F. (Round your answer to three decimal places.) y ≈

OpenStudy (anonymous):

x^2=pi/2,x^2=5*pi/2;